# Beyond Scores: A Modular RAG-Based System for Automatic Short Answer Scoring with Feedback

### v1

@@ V-B2 Human Evaluation
To ensure the quality of the feedback generated by our system, we conducted a human evaluation with five expert teachers in Information Technology and Networks. The raters assessed 108 randomly selected samples from 12 models, focusing on accuracy and clarity. Feedback was rated on a 5-point Likert scale ranging from -2 to 2, where -2 indicated strongly inaccurate or unclear feedback, 0 indicated neutral feedback, and 2 indicated strongly accurate or clear feedback. Figures 6a and 6b illustrate the human evaluation results for samples with existing student answers. Table V presents the overall results, excluding no-response samples.
For the baseline models, human evaluation was not conducted. Table VI provides a sample of the feedback generated by the baseline model and samples from our ASAS-F system. The feedback generated by the baseline model is repetitive and lacks the depth needed for comprehensive understanding. A high degree of similarity could artificially inflate the statistical evaluation metrics like BLEU scores. The high overlap between training inputs and generated outputs signals that the model may not truly be learning to adapt to novel inputs but is instead reproducing familiar patterns. Thus, while the baseline system might achieve high scores on automated metrics due to repetition of training data, its semantic contribution may be minimal compared to the more informative and contextually appropriate feedback generated by our system. More examples of comparisons between the feedback generated by the baseline models and our ASAS-F system can be found in Appendix B.
@@
In the few-shot approach, LLMs often mimicked the style of the provided reference answers or examples as shown in the first example in Figure 7. Many raters commented the feedback that was too brief, lacking sufficient context or explanation. Although the feedback was factually correct, it did not provide the depth needed for comprehensive understanding, which could be critical in educational settings. Conversely, in the zero-shot approach, the generated feedback was often longer and appeared more detailed. However, as the second example in Figure 7 demonstrates, while some of this feedback was labeled as accurate, it was actually incorrect when compared to the reference answers. This highlights the challenge of calibrating LLMs, as they may present incorrect feedback with high confidence [39], making it difficult to detect when they are hallucinating [40], especially when evaluating the feedback in isolation.
VI Discussion
@@ VII Conclusion
Acknowledgment
@@ Appendix B Output Generation Examples
This section provides a comparison of feedback samples generated by the baseline models and our system. Tables VII, VIII, and IX present feedback generated by the baseline models and samples from our system for the same answers. These samples show clear instances of phrase reuse from the fine-tuning process from the baseline, highlighting limited generalization capabilities compared to the more informative and contextually appropriate feedback generated by our system.
@@ Appendix C Typed Predictor Error Analysis
Table X presents the error rates for each model. Although typed predictors adhere to strict standards, some errors are unavoidable due to the complexity of the input data and the constraints on the output structure. On average, the Llama models exhibited lower error rates compared to the Mistral models, suggesting that the Llama models were better at adhering to the predefined type constraints. Among the Llama models, the smaller Llama3:8b achieved the lowest error rate at 2.04%, followed by the Llama3:70b with a rate of 3.0%. In contrast, the Mistral:7b model had an error rate of 4.16%, while the Mixtral:8x22b model recorded the highest average error rate at 7.32%.

### v2

@@ 5.2.2 Human Evaluation
To ensure the quality of the feedback generated by our system, we conducted a human evaluation with five expert teachers in Information Technology and Networks. The raters assessed 108 randomly selected samples from 12 models, focusing on accuracy and clarity. Feedback was rated on a 5-point Likert scale ranging from -2 to 2, where -2 indicated strongly inaccurate or unclear feedback, 0 indicated neutral feedback, and 2 indicated strongly accurate or clear feedback. Figures 6a and 6b illustrate the human evaluation results for samples with existing student answers. Table 5 presents the overall results, excluding no-response samples.
For the baseline models, human evaluation was not conducted. Table 6 provides a sample of the feedback generated by the baseline model and samples from our ASAS-F system. The feedback generated by the baseline model is repetitive and lacks the depth needed for comprehensive understanding. A high degree of similarity could artificially inflate the statistical evaluation metrics like BLEU scores. The high overlap between training inputs and generated outputs signals that the model may not truly be learning to adapt to novel inputs but is instead reproducing familiar patterns. Thus, while the baseline system might achieve high scores on automated metrics due to repetition of training data, its semantic contribution may be minimal compared to the more informative and contextually appropriate feedback generated by our system. More examples of comparisons between the feedback generated by the baseline models and our ASAS-F system can be found in Appendix B.
@@
In the few-shot approach, LLMs often mimicked the style of the provided reference answers or examples as shown in the first example in Figure 7. Many raters commented the feedback that was too brief, lacking sufficient context or explanation. Although the feedback was factually correct, it did not provide the depth needed for comprehensive understanding, which could be critical in educational settings. Conversely, in the zero-shot approach, the generated feedback was often longer and appeared more detailed. However, as the second example in Figure 7 demonstrates, while some of this feedback was labeled as accurate, it was actually incorrect when compared to the reference answers. This highlights the challenge of calibrating LLMs, as they may present incorrect feedback with high confidence Zhou et al. (2024), making it difficult to detect when they are hallucinating Ji et al. (2023), especially when evaluating the feedback in isolation.
6 Discussion
@@ 7 Conclusion
Acknowledgements
This work was partially supported by JST (Japan Science and Technology Agency), the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2132, and JSPS (Japan Society for the Promotion of Science) KAKENHI No. JP21H00907 and JP23H03511.
@@ Appendix B Output Generation Examples
This section provides a comparison of feedback samples generated by the baseline models and our system. Tables 7, 8, and 9 present feedback generated by the baseline models and samples from our system for the same answers. These samples show clear instances of phrase reuse from the fine-tuning process from the baseline, highlighting limited generalization capabilities compared to the more informative and contextually appropriate feedback generated by our system.
@@ Appendix C Typed Predictor Error Analysis
Table 10 presents the error rates for each model. Although typed predictors adhere to strict standards, some errors are unavoidable due to the complexity of the input data and the constraints on the output structure. On average, the Llama models exhibited lower error rates compared to the Mistral models, suggesting that the Llama models were better at adhering to the predefined type constraints. Among the Llama models, the smaller Llama3:8b achieved the lowest error rate at 2.04%, followed by the Llama3:70b with a rate of 3.0%. In contrast, the Mistral:7b model had an error rate of 4.16%, while the Mixtral:8x22b model recorded the highest average error rate at 7.32%.